# SoK: Evaluating Jailbreak Guardrails for Large Language Models

### v1

@@ 5.2 Benchmark Evaluation
Shifting focus to PGR, presented in Table III, GuardReasoner (Post) achieves the best PGR of 0.141. Despite its superior precision in identifying malicious inputs, GuardReasoner (Post) does not attain state-of-the-art (SOTA) overall defense performance. A plausible explanation is its potential operational overlap with the target LLM’s intrinsic safety mechanisms. That is, there might be a significant number of instances where GuardReasoner (Post) identifies a response as safe, and concurrently, the target LLM also recognizes the harmful nature of the query and refuses to respond, thereby diminishing the unique contribution of GuardReasoner (Post) to the ASR reduction when compared to a guardrail like GuardReasoner (Pre) which operates on a different paradigm.
@@
Efficiency. The efficiency of guardrails is a critical factor for practical deployment, which we assess in terms of latency and GPU memory consumption. Figure 2(a) illustrates the extra delay introduced by different guardrail methodologies when processing normal inputs from the AlpacaEval and OR-Bench datasets. Perplexity Filter, Llama Guard, SelfDefend and PromptGuard stand out with the negligible latency. In contrast, GuardReasoner (Pre) and GradientCuff impose the most significant delays, with GuardReasoner (Pre) being particularly notable. This suggests that the profound reasoning capabilities that afford GuardReasoner (Pre) its enhanced defense performance come at the cost of increased processing time. The majority of other guardrails maintain an additional delay generally not exceeding 0.5 seconds.
@@
From the perspective of GPU memory utilization, depicted in Figure 2(b), GuardReasoner (Pre) again registers the highest memory footprint, consistent with its complex reasoning architecture. Conversely, SmoothLLM, GradientCuff, and PromptGuard are the most memory-efficient, with their consumption approaching negligible levels. This highlights a clear trade-off between the sophistication of the defense mechanism and its resource intensiveness.
@@
Utility. Beyond security and efficiency, the utility of a guardrail, specifically its ability to not impede benign user interactions, is paramount. We measure the FPR on AlpacaEval and OR-Bench datasets, as shown in Figure 2(c). A higher FPR indicates a greater propensity to incorrectly flag legitimate prompts as malicious. SelfDefend (Direct) exhibits the highest FPR on OR-Bench, at 0.221. On AlpacaEval, GradientCuff records the highest FPR of 0.083. These figures suggest that these guardrails have a higher likelihood of intercepting normal user queries. Other guardrails with comparatively high FPRs include SmoothLLM, SelfDefend (Intent), WildGuard (Pre), and GuardReasoner (Pre). Although these three methods demonstrate strong defense performance (low ASR), their elevated FPRs underscore a critical trade-off between security and utility. Systems that are highly stringent in blocking threats may inadvertently penalize legitimate interactions, diminishing the overall user experience.

### v2

@@ 5.2 Benchmark Evaluation
Shifting focus to PGR, presented in Table III, GuardReasoner (Post) achieves the best PGR of 0.141. Despite its superior precision in identifying malicious inputs, GuardReasoner (Post) does not attain state-of-the-art (SOTA) overall defense performance. A plausible explanation is its potential operational overlap with the target LLM’s intrinsic safety mechanisms. That is, there might be a significant number of instances where GuardReasoner (Post) identifies a response as safe, and concurrently, the target LLM also recognizes the harmful nature of the query and refuses to respond. Therefore, this overlap diminishes the unique contribution of GuardReasoner (Post) to the ASR reduction when compared to a guardrail like GuardReasoner (Pre) which operates on a different paradigm.
@@
Efficiency. The efficiency of guardrails is a critical factor for practical deployment, which we assess in terms of latency and GPU memory consumption. Figure 2(a) illustrates the extra delay introduced by different guardrail methodologies when processing normal inputs from the AlpacaEval and OR-Bench datasets. Perplexity Filter, Llama Guard, SelfDefend and PromptGuard stand out with negligible latency. In contrast, GuardReasoner and GradientCuff impose the most significant delays, with GuardReasoner (Post) being particularly notable. This suggests that the profound reasoning capabilities that afford GuardReasoner its enhanced defense performance come at the cost of increased processing time. The majority of other guardrails maintain an additional delay generally not exceeding 0.5 seconds.
@@
From the perspective of GPU memory utilization, depicted in Figure 2(b), GuardReasoner again registers the highest memory footprint, consistent with its complex reasoning architecture. Conversely, SmoothLLM, GradientCuff, and PromptGuard are the most memory-efficient, with their consumption approaching negligible levels. This highlights a clear trade-off between the sophistication of the defense mechanism and its resource intensiveness.
@@
Utility. Beyond security and efficiency, the utility of a guardrail, specifically its ability not to impede benign user interactions, is paramount. We measure the FPR on AlpacaEval and OR-Bench datasets, as shown in Figure 2(c). A higher FPR indicates a greater propensity to incorrectly flag legitimate prompts as malicious. SelfDefend (Direct) exhibits the highest FPR on OR-Bench, at 0.221. On AlpacaEval, GradientCuff records the highest FPR of 0.083. These figures suggest that these guardrails have a higher likelihood of intercepting normal user queries. Other guardrails with comparatively high FPRs include SmoothLLM, SelfDefend (Intent), WildGuard (Pre), and GuardReasoner (Pre). Although these three methods demonstrate strong defense performance (low ASR), their elevated FPRs underscore a critical trade-off between security and utility. Systems that are highly stringent in blocking threats may inadvertently penalize legitimate interactions, diminishing the overall user experience.
Figure 3: The heatmap of guardrails.
@@ 6 Practical Insights & Implications
7 Discussion
7.1 Limitations
While this work provides a comprehensive analysis of LLM jailbreak guardrails, several limitations should be acknowledged. First, our efficiency evaluations were conducted on specific hardware configurations (NVIDIA H800 GPUs), and the absolute latency and memory consumption metrics may vary across different hardware platforms and optimization techniques such as quantization. However, the relative performance trends and trade-offs among guardrails are expected to remain consistent.
Besides, our evaluation relies on LLM-as-a-judge (GPT-4o) for assessing jailbreak success, which, while efficient for large-scale evaluation, may have inherent limitations in reliability. To validate this approach, we hired three Ph.D. students to evaluate the jailbreak success of 100 randomly selected attack samples. Their average agreement with the LLM’s judgments was 0.9533, and their Fleiss’ Kappa was 0.9319, indicating strong alignment. Nevertheless, future work could benefit from more robust evaluation methodologies that mitigate potential biases in automated assessment.
7.2 Future Work
Based on our analysis, we identify several promising directions for future research in LLM jailbreak guardrails:
  • Cross-Attack Robust Guardrails: Current guardrails demonstrate limited generalization across different attack modalities. Future work should develop unified defense mechanisms that can effectively handle diverse threats, including jailbreaks, prompt injections, and other adversarial manipulations simultaneously. This requires guardrails that can recognize attack patterns beyond jailbreaks.
  • Adaptive and Self-Evolving Defenses: As jailbreak techniques continuously evolve, fixed guardrails face the risk of rapid obsolescence. Research is needed on self-adaptive guardrails that can learn from new attack patterns and update their detection capabilities without complete retraining. This could include online learning approaches, anomaly detection in deployment environments, and automated defense refinement.
  • Multimodal and Cross-Modal Guardrails: With the rise of multimodal LLMs, new vulnerabilities emerge at the intersection of different modalities. Future research should address the unique challenges of multimodal jailbreaks [107] and develop guardrails that can analyze and protect across text, image, audio, and video inputs while maintaining efficiency.
@@ 8 Conclusion
Acknowledgments
We thank the reviewers and the shepherd for their constructive comments. The HKUST authors are supported in part by an RGC CRF grant under contract C6015- 23G and a research fund provided by HSBC. Daoyuan Wu was partially supported by Lingnan Grant SUG-002/2526.
Appendix A Evaluation Results on Vicuna-13b-v1.5
We extend our comprehensive evaluation to another widely-used open-source model, Vicuna-13b-v1.5, to assess the generalization of different guardrails. The detailed results are presented in Table IV.
First, a salient observation is that Vicuna-13b-v1.5 is considerably more susceptible to jailbreak attacks compared to Llama-3. This increased vulnerability is evident from the substantially higher Attack Success Rates (ASR) across almost all attack categories, indicating a weaker inherent safety alignment in Vicuna.
Second, we note a significant performance degradation for certain defenses when applied to Vicuna-13b-v1.5. For example, the efficacy of GradSafe and GradientCuff diminishes. GradientCuff, which showed marked effectiveness against the X-Teaming multi-turn attack on Llama-3, fails to maintain this advantage on Vicuna-13b-v1.5. This decline can be attributed to their nature as intra-processing guardrails, which heavily rely on the internal representations and alignment of the target LLM. Consequently, a less well-aligned model like Vicuna-13b-v1.5 compromises their defensive mechanism.
Despite these differences, we also observe consistent performance patterns. GuardReasoner (Pre) and GuardReasoner (Post) continue to exhibit state-of-the-art defense capabilities. GuardReasoner (Pre) achieves the best overall ASR of 0.156, while GuardReasoner (Post) records the best overall PGR of 0.192. This sustained excellence underscores that the robust defense mechanism of GuardReasoner is largely independent of the target LLM, positioning it as a more universally applicable and reliable guardrail.
TABLE V: The ASR ($\downarrow$) / PGR ($\downarrow$) results for the target LLM (GPT-4-0125-Preview).
GPT-4-0125-Preview 0.192/- 0.220/- 0.710/- 0.340/- 0.750/- 0.220/- 0.029/- 0.200/- 0.940/- 0.400/-
PerplexityFilter 0.192/1.000 0.000/0.060 0.710/1.000 0.260/1.000 0.720/1.000 0.220/1.000 0.029/1.000 0.200/1.000 0.930/1.000 0.362/0.896
SmoothLLM 0.185/0.365 0.150/0.270 0.480/0.850 0.540/0.630 0.760/0.940 0.300/0.780 0.054/1.000 0.192/0.905 0.980/0.980 0.405/0.747
Llama Guard (Pre) 0.100/0.563 0.130/0.460 0.540/0.770 0.250/0.510 0.660/0.840 0.220/0.850 0.029/0.952 0.200/0.998 0.940/0.990 0.341/0.770
Llama Guard (Post) 0.110/0.110 0.180/0.180 0.530/0.530 0.250/0.250 0.720/0.720 0.220/0.220 0.029/0.029 0.200/0.200 0.950/0.950 0.354/0.354
SelfDefend (Direct) 0.031/0.253 0.060/0.100 0.160/0.230 0.100/0.210 0.110/0.160 0.070/0.350 0.019/0.740 0.102/0.693 0.870/0.910 0.169/0.405
SelfDefend (Intent) 0.039/0.310 0.040/0.090 0.080/0.160 0.100/0.190 0.100/0.180 0.060/0.140 0.025/0.568 0.128/0.768 0.910/0.930 0.165/0.371
WildGuard (Pre) 0.004/0.033 0.000/0.020 0.020/0.020 0.030/0.060 0.010/0.010 0.140/0.530 0.029/0.797 0.127/0.768 0.810/0.880 0.130/0.346
WildGuard (Post) 0.035/0.035 0.030/0.030 0.070/0.070 0.060/0.060 0.180/0.180 0.140/0.140 0.029/0.029 0.142/0.142 0.920/0.920 0.178/0.178
Prompt Guard 0.000/0.000 0.020/0.050 0.410/0.630 0.310/0.960 0.000/0.000 0.220/0.970 0.029/1.000 0.195/0.993 0.910/1.000 0.233/0.623
GuardReasoner (Pre) 0.000/0.009 0.000/0.000 0.030/0.030 0.030/0.040 0.050/0.070 0.080/0.290 0.016/0.349 0.113/0.740 0.890/0.960 0.134/0.276
GuardReasoner (Post) 0.021/0.021 0.030/0.030 0.040/0.040 0.080/0.080 0.240/0.240 0.100/0.100 0.029/0.029 0.110/0.110 0.930/0.930 0.176/0.176
Appendix B Evaluation Results on GPT-4-0125-Preview
To further validate the generalizability of our findings, we evaluate guardrail performance on GPT-4-0125-Preview, a representative closed-source model. The comprehensive results are presented in Table V, Figure 6, 7, and 8.
Table V demonstrates that the defense performance metrics (ASR/PGR) on GPT-4 consistently align with those observed in Llama-based evaluations. Notably, PerplexityFilter achieves the highest average PGR, while GuardReasoner (Post) maintains the lowest. Furthermore, the multi-turn X-Teaming attack continues to exhibit exceptionally high ASR/PGR across all guardrails, reinforcing the findings from RQ3 regarding session-level defense limitations.
Figure 6 reveals consistent efficiency patterns: GuardReasoner incurs substantial latency due to its reasoning overhead, and post-processing guardrails against jailbreaks universally introduce greater delay than their pre-processing counterparts, directly addressing RQ1. The pronounced latency of SmoothLLM stems from its defense mechanism requiring 10 perturbed prompt copies. This issue is further exacerbated by inefficient asynchronous API access to GPT-4 when compared to batched local inference.
Figure 7 illustrates the GPU memory usage of different guardrails when applied to GPT-4. These GPU memory consumptions (in GB) for GPT-4 align with the conclusions drawn from Llama-3. Specifically: (1) GuardReasoner exhibits the highest memory footprint. (2) The memory overhead for SmoothLLM and PromptGuard is negligible. (3) LLM-based guardrails such as Llama Guard, SelfDefend, WildGuard, and GuardReasoner consume more GPU memory compared to rule-based guardrails (e.g., SmoothLLM). These have the same answer to RQ2.
The utility results (FPR) in Figure 8 for GPT-4 corroborate the conclusions drawn from Llama-3. Specifically: (1) SelfDefend (Direct) exhibits the highest FPR on OR-Bench. (2) Token-level guardrails (SmoothLLM, SelfDefend (Direct)) demonstrate relatively pronounced FPRs. (3) Session-level guardrails (those with a “Post” suffix) consistently achieve markedly lower FPRs than their sequence-level counterparts (those with a “Pre” suffix). These findings consistently address RQ4.
These results demonstrate that our SEU framework and taxonomic insights generalize effectively to large-scale black-box models, affirming the robustness of our findings.
Appendix C Meta-Review
The following meta-review was prepared by the program committee for the 2026 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
C.1 Summary
This paper presents a defense-oriented systematization of knowledge on protecting large language models against jailbreak attacks. It proposes a taxonomy of jailbreak techniques and categorizes guardrail defenses along several key dimensions: the underlying technical paradigm, the granularity of protection, the degree of reactivity, their applicability across contexts, and the level of explainability they provide. In addition, the paper includes an evaluation of existing defenses, offering comparative insights into their effectiveness.
C.2 Scientific Contributions
  • A systematization of defenses for LLMs
C.3 Reasons for Acceptance
  1. 1.
    This paper proposes a comprehensive taxonomy of LLM guardrails. The presentation is clear and has good coverage of existing work.
  2. 2.
    The paper also proposes a new evaluation framework that considers real-world trade-offs like latency, computational cost and usability.
  3. 3.
    The evaluation framework is demonstrated with various defenses and LLM models. This offers concrete benchmarks for comparing defenses.